OpenStudy (anonymous):

A puppy has a mass of 4,500 grams and a kitten has a mass of 450,000 milligrams.which statement is true about the animals? A. The puppy has a greater mass B.the kitten has the greater mass C.they have the same mass. @abyss @31356 @mathstudent55

OpenStudy (anonymous):

a

OpenStudy (anonymous):

1000mg = 1gm

OpenStudy (mathstudent55):

Let's convert both to the same unit and compare. SInce the mass of the puppy is already in grams, let's convert the mass of a kitten to grams.

OpenStudy (mathstudent55):

1 g = 1000 mg That means \( \dfrac{1~g}{1000~mg} = \dfrac{1000~mg}{1~g} =1\) Since 450,000 mg has mg in the numnerator, and we want to cacel it, we multiply by the conversion factor that has mg in the denominator. \(450,000~mg \times \dfrac{1~g}{1000~mg} = 450 ~g\)

OpenStudy (mathstudent55):

Now you see that the puppy has a mass of 4,500 g, and a kitten has a mass of 450 g. Now it's easy to compare them.